OpenStudy (anonymous):

Given f(x) and g(x), write g(f(x)). f(x)=3x−4;g(x)=√x+ 2

OpenStudy (anonymous):

f(x)=3x−4;g(x)=√x+ 2 \[=g(f(x))\] \[=g(3x-4)\] \[=\sqrt{3x-4}+2\]

OpenStudy (anonymous):

Get the picture?

OpenStudy (anonymous):

Thank you guys so much!!